Grade 12Algebra

The vapour pressure of pure water at 30°C is 31.80 mm of hg. How many grams of urea(molecular mass =60) should be dissolved in 1000 g of water to lower the vapour pressure by 0.25 mm of hg?

Manan patidar
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
As we know
 
p° - p /p° = W2 M1 /M2 W1
 
Now put the values
 
0.25 /31.8 = W2 *18 /(60* 100)
 
On calculation,
 
W2 = 2.62 gm